# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v3

@@ Abstract
Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models
Anonymous ACL submission
@@ 1 Introduction
Personal values are broad, fundamental motivations behind individual and collective actions. They serve as guiding principles that shape perceptions, cognition, and behaviors across situations and over time. As Large Language Models (LLMs) continue to permeate and transform human society, researchers have increasingly sought to evaluate, understand, and align LLM values (ye2025gpv; ren2024valuebench; rao2024normad; meadows2024localvaluebench; kovavc2024stick; jiang2024raising; rozen2024llms; yao2024clave; zhang2023measuring; sorensen2024value; moore2024large; demszky2023using). Rather than relying solely on specific risk metrics, implicit preference modeling, or atomic ethical principles, framing these studies through the lens of intrinsic values schwartz2012overview offers a more comprehensive, adaptable, and transparent approach yao2023value_fulcra; biedma2024beyond; yao2023alignment_goals.
@@
However, value systems used in existing research—primarily Schwartz’s values—are designed for humans. A psychologically informed value system specifically tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems biedma2024beyond has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in § 2.3 and Appendix C.
@@
To address this gap, we introduce a generative psycho-lexical approach (GPLA), a novel methodology that leverages LLMs to construct a value system grounded in psychological principles. This approach is based on the psycho-lexical hypothesis, which suggests that all salient values are captured in language. GPLA adopts an agentic framework and involves five sequential steps: 1) extracting perceptions from the corpus, 2) identifying values behind perceptions, 3) filtering values, 4) non-reactive value measurements of test subjects, and 5) structuring the value system. In contrast to traditional methods that manually compile value lexicons and design questionnaires to collect self-report data aavik2002structure; crectu2012psycho; Klages1929-KLATSO-5; saucier2000isms, GPLA automates the entire process and features two key advantages.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§ B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on forced-choice self-report are demonstrated to be unreliable and biased for humans ponizovskiy2020development; bardi2008new and AI ye2025gpv; rottger2024political; dominguez2023questioning. Instead, GPLA measures values through analyzing unstructured responses to arbitrary prompts, which mitigates response biases and enables scalable and self-adaptive value measurement ye2025gpv.
@@
In applying GPLA to 33 LLMs, each steered with 21 different value-anchoring prompts rozen2024llms, we develop a five-factor value system encompassing Social Responsibility, Risk-taking, Rule-following, Self-competence, and Rationality. We demonstrate the system’s reliability and explore its theoretical and practical implications. To benchmark LLM value systems, we propose three tasks based on psychometric principles devon2007psychometric and AI priorities ji2023ai_alignment; anwar2024foundational: Confirmatory Factor Analysis for assessing structural validity, LLM Safety Prediction for examining predictive validity, and LLM Value Alignment for evaluating representational power. These tasks confirm the validity and utility of our system in evaluating, understanding, and aligning the intrinsic values of LLMs, when compared to canonical Schwartz’s values schwartz2012overview.
@@
We summarize our contributions as follows. 1) We introduce the Generative Psycho-lexical Approach (GPLA), a novel LLM-driven and data-driven methodology to construct psychologically grounded value systems. 2) With GPLA, we construct a five-factor value system for LLMs. We demonstrate its reliability and explore its theoretical and practical implications. 3) We present three benchmarking tasks for LLM value systems, and establish the superior validity and utility of the proposed system.
@@ 2.1 Values, Value Measurement, and Theory-Driven Value Systems
Values are broad, fundamental beliefs or principles that guide individuals’ behaviors, cognition, and decisions schwartz1987toward. They represent what people consider important in life and serve as motivational forces when evaluating actions, people, and events schwartz2012overview. Unlike other psychological constructs such as personalities, values are stable, motivational, and transituational sagiv2022personal. These characteristics make values uniquely powerful for understanding human behaviors bardi2003values, evaluating individual and cultural differences schwartz2006theory, and resolving conflicts rooted in clashing priorities sen1986social.
@@
Values are measurable constructs schwartz2001extending. Value measurement seeks a quantitative assessment of the importance individuals or organizations assign to specific values. Traditionally, it relies on tools like self-report questionnaires schwartz2001extending; maio2010mental, behavioral observation lonnqvist2013personal; maio2009changing; fischer2011whence, and experimental techniques sagiv2011compete; yamagishi2013behavioral; murphy2014social. However, these tools are hindered by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data ponizovskiy2020development; boyd2015values; bardi2008new. Recently proposed Generative Psychometrics for Values (GPV) ye2025gpv leverage LLMs to dynamically parse unstructured texts into perceptions and measure values therein. GPV is shown to be more scalable, reliable, and flexible than traditional tools in measuring LLM values, and is utilized for non-reactive value measurement in this work.
@@
Isolated values are structured into a value system for a holistic understanding of their relationships and implications schwartz2012overview. Theory-driven value systems like Schwartz’s theory of basic human values schwartz1992universals and functional theory of values gouveia2014functional are rooted in theoretical hypotheses and preconceptions of psychologists. These systems are established prior to data collection and analysis, and are later validated through empirical studies. However, theory-driven approaches are constrained by the subjectivity of the theorists, limited coverage, and an inability to adapt to evolving values or specific contexts ponizovskiy2020development; raad2017reply.
@@ 2.2 Psycho-Lexical Approach
In contrast to the theory-driven approaches, the psycho-lexical approach organizes psychological constructs such as values, personality traits, and social attitudes aavik2002structure; crectu2012psycho; Klages1929-KLATSO-5; saucier2000isms, under a data-driven paradigm. It operates on the fundamental principle that language naturally evolves to capture salient and socially relevant individual differences de2016values. The pioneering attempts date back to the early 20th century allport1936trait; galton1950character, and the paradigm has then been developed and extended to different psychological constructs for decades cattell1957personality; john1988lexical; aavik2002structure; ponizovskiy2020development; mai2023exploring; garrashi2024personality.
Most of these works, taking values as an example, involve the following steps: 1) compilation of value-descriptive terms, mostly from dictionaries and thesauruses; 2) refinement and reduction to eliminate synonyms, ambiguous terms, and those that are not commonly used or understood; 3) value measurement, through collecting self-reports and peer ratings, to identify underlying correlations between the value descriptors; and 4) uncovering the hidden value factors or dimensions through statistical methods like principal component analysis or factor analysis.
In this work, we aim to harness the extensive knowledge and semantic understanding of LLMs to address the limitations of traditional psycho-lexical approaches. These limitations include: 1) extensive manual labor required to compile and refine the lexicons, 2) insufficient coverage of values of different linguistic forms, 3) lack of criteria for prioritizing values lexicons when filtering, 4) the responses bias and inscalability of self-report value measurement, and 5) the inability to adapt to changing values or specific contexts.
@@ 2.3 From Human Values to LLM Values
The rise of LLMs also brings the critical need to evaluate, understand, and align their values. Extensive works focus on evaluating the values of LLMs using traditional self-report tools li2022gpt; li2024evaluatingpsychologicalsafetylarge; huang2024humanity; safdari2023-personality-traits-in-llm; safdari2023personality, static customized inventories ren2024valuebench; meadows2024localvaluebench; jiang2024evaluating, dynamically generated inventories jiang2024raising, or directly from the model’s free-form outputs ye2025gpv; yao2024clave; yao2023value_fulcra; yao2025leaderboard. Other works attempt to understand the LLM values in aspects like the consistency of shown values rozen2024llms; moore2024large; rottger2024political, the ability to reason about human values ganesan2023systematic; sorensen2024value; jiang2024can; strachan2024testing, and the ability to express or role-play human values jiang-etal-2024-personallm; zhang2023measuring; kang2024causal; zhang2025extrapolating; li2024quantifying; huang2024social. Another line of research aligns LLM values with human values, setting the alignment goal as specific risk metrics lin2024towards; gunjal2024detecting, human demonstrations dubois2024alpacafarm; kopf2024openassistant; alpaca, implicit preference modeling ouyang2022training; rafailov2024dpo; zhong2024panacea, or intrinsic values bai2022hh; yao2023value_fulcra; bai2022constitutional. Among different alignment goals, intrinsic values, as transituational decision-guiding principles, demonstrate unique advantages yao2023alignment_goals. However, prior works are mostly based on Schwartz’s value system, which is established and validated for humans and may not necessarily capture the unique LLM psychology. Therefore, it is essential to construct a value system for LLMs grounded in psychological theories and principles, based on which we can more effectively evaluate, understand, and align the values of LLMs.
One preliminary attempt biedma2024beyond of constructing an LLM value system faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is prone to response bias and lacks comprehensive coverage; 2) the derivation of value correlations is not theoretically grounded and lacks empirical validity; and 3) the resulting value system has not been evaluated in terms of its reliability, validity, or utility. Appendix C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
This section presents the Generative Psycho-Lexical Approach (GPLA), illustrated schematically in Fig. 1 and algorithmically in Algorithm 1, as the foundation for constructing the LLM value system. GPLA adopts an agentic framework, comprising three LLM agents and four sequential steps, to systematically collect, filter, measure, and structure the LLM value system.
@@
In line with the established psycho-lexical approach de2016values, we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs from ValueBench ren2024valuebench, GPV ye2025gpv, BeaverTails ji2024beavertails, and ValueLex biedma2024beyond. These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in § A.1.
@@
Then, we extract perceptions from the collected corpus using the Perception Parser . Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools ye2025gpv. These perceptions are then mapped to underlying values through the Value Generator , with their frequency of occurrence recorded. We instantiate following GPV ye2025gpv and using Kaleido model sorensen2024value. In § B.1, we confirm the comprehensive coverage of the generated values.
@@
To ensure the value lexicons are concise and representative, we filter out duplicates using Rouge scores and embedding similarity following the validated setup in sorensen2024value. When two values exhibit significant semantic overlap, we retain the one more frequently observed, as higher frequency signifies greater priority ponizovskiy2020development. This process results in value lexicons that minimize redundancy while preserving semantic coverage.
@@
Structuring the LLM value system requires identifying correlations between values derived from LLM value measurements. To this end, we adopt GPV ye2025gpv, instantiated with and , to measure the value orientations of LLM subjects on values . GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein. We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (§ A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system is then constructed through Principal Component Analysis (PCA) ponizovskiy2020development, which uncovers the underlying value factors and hierarchical relationships (loadings of on ).
4 Validation and Advantages of GPLA
4.1 Are LLMs Capable for GPLA?
LLMs assume three roles in GPLA: Perception Parser, Value Generator, and Value Evaluator. This section validates the capabilities of LLMs in these roles.

### v4

@@ Abstract
Generative Psycho-Lexical Approach for Constructing
Value Systems in Large Language Models
Haoran Ye††thanks: Equal contribution.,1, Tianze Zhang11footnotemark: 1,1,2, Yuhang Xie11footnotemark: 1,1, Liyuan Zhang1, Yuanyi Ren1, Xin Zhang3,4, Guojie Song††thanks: Corresponding author.,1,5 1State Key Laboratory of General Artificial Intelligence, School of Intelligence Science and Technology, Peking University 2Yuanpei College, Peking University 3School of Psychological and Cognitive Sciences, Peking University 4Key Laboratory of Machine Perception (Ministry of Education), Peking University 5PKU-Wuhan Institute for Artificial Intelligence {hrye, ericzhang, yuhangxie, zly2003}@stu.pku.edu.cn {yyren, zhang.x, gjsong}@pku.edu.cn
@@ 1 Introduction
Personal values are broad, fundamental motivations behind individual and collective actions. They serve as guiding principles that shape perceptions, cognition, and behaviors across situations and over time. As Large Language Models (LLMs) continue to permeate and transform human society, researchers have increasingly sought to evaluate, understand, and align LLM values (Ye et al., 2025b; Ren et al., 2024; Rao et al., 2024; Meadows et al., 2024; Kovač et al., 2024; Jiang et al., 2024b; Rozen et al., 2024; Yao et al., 2024b; Sorensen et al., 2024a). Rather than relying solely on specific risk metrics, implicit preference modeling, or atomic ethical principles, framing these studies through the lens of intrinsic values Schwartz (2012) offers a more comprehensive, adaptable, and transparent approach Yao et al. (2024a); Biedma et al. (2024); Yao et al. (2023).
@@
However, value systems used in existing research—primarily Schwartz’s values—are established for humans. A psychologically informed value system tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems Biedma et al. (2024) has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in § 2.3 and Appendix C.
@@
To address this gap, we introduce a generative psycho-lexical approach (GPLA), a novel methodology that leverages LLMs to construct a value system grounded in psychological principles. This approach is based on the psycho-lexical hypothesis, which suggests that all salient values are captured in language. GPLA adopts an agentic framework and involves five sequential steps: 1) extracting perceptions from the corpus, 2) identifying values behind perceptions, 3) filtering values, 4) non-reactive value measurements of test subjects, and 5) structuring the value system. In contrast to traditional methods that manually compile value lexicons and design questionnaires to collect self-report data Aavik and Allik (2002); Creţu et al. (2012); Saucier (2000), GPLA automates the entire process and features two key advantages.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§ B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on self-report are demonstrated to be unreliable and biased for humans Ponizovskiy et al. (2020) and AI Dominguez-Olmedo et al. (2023); Röttger et al. (2024); Ye et al. (2025b). Instead, GPLA measures values within unstructured responses to arbitrary prompts, which mitigates response biases and enhances scalability and validity Ye et al. (2025b).
@@
In applying GPLA to 33 LLMs, each steered with 21 different value-anchoring prompts Rozen et al. (2024), we develop a five-factor value system encompassing Social Responsibility, Risk-taking, Rule-following, Self-competence, and Rationality. We demonstrate the system’s reliability and explore its theoretical and practical implications. To benchmark LLM value systems, we propose three tasks based on psychometric principles DeVon et al. (2007) and AI priorities Anwar et al. (2024): Confirmatory Factor Analysis for assessing structural validity, LLM Safety Prediction for examining predictive validity, and LLM Value Alignment for evaluating representational power. These tasks confirm the validity and utility of our system in evaluating, understanding, and aligning the intrinsic values of LLMs, when compared to canonical Schwartz’s values Schwartz (2012).
@@
In summary, we contribute: 1) the Generative Psycho-Lexical Approach (GPLA), a novel LLM- and data-driven methodology for constructing psychologically grounded value systems; 2) a five-factor value system constructed using GPLA for LLMs, with demonstrated reliability and an exploration of its theoretical and practical implications; and 3) three benchmarking tasks for LLM value systems, through which we establish the superior validity and utility of the proposed system.
@@ 2.1 Values, Value Measurement, and Theory-Driven Value Systems
Values are broad, fundamental beliefs and guiding principles, serving as motivational forces when evaluating actions, people, and events Schwartz (2012). Unlike other psychological constructs such as personalities, values are stable, motivational, and transituational Sagiv and Schwartz (2022), rendering them uniquely powerful across disciplines for investigating decision-making and social dynamics Sen (1986); Schwartz (2006).
@@
Values are measurable constructs Schwartz et al. (2001). Value measurement seeks a quantitative assessment of the importance an agent assigns to specific values. Traditionally, it relies on tools like self-report questionnaires Maio (2010), behavioral observation Lönnqvist et al. (2013), and experimental techniques Murphy and Ackermann (2014), which are hindered by response biases, resource demands, and inabilities to capture authentic behaviors Ponizovskiy et al. (2020); Boyd et al. (2015). Recently proposed Generative Psychometrics for Values (GPV) Ye et al. (2025b) leverage LLMs to dynamically parse unstructured texts into perceptions and measure values therein. GPV is shown to be more scalable, valid, and flexible than traditional tools in measuring LLM values, and is utilized for non-reactive value measurement in this work.
@@
Isolated values are structured into a value system for a holistic understanding of their hierarchy and interplay Schwartz (2012). Theory-driven value systems like Schwartz’s theory of basic human values Schwartz (1992) and functional theory of values Gouveia et al. (2014) are rooted in theoretical hypotheses and preconceptions of psychologists. These systems are established prior to data collection and analysis, and are later validated through empirical studies. However, theory-driven approaches are constrained by the subjectivity of the theorists, limited coverage, and inadaptability to evolving values or specific contexts Ponizovskiy et al. (2020); Raad et al. (2017).
@@ 2.2 Psycho-Lexical Approach
In contrast to the theory-driven approaches, the psycho-lexical approach organizes psychological constructs such as values, personality traits, and social attitudes Aavik and Allik (2002); Creţu et al. (2012); Klages and Johnson (1929); Saucier (2000), under a data-driven paradigm. It operates on the fundamental principle that language naturally evolves to capture salient and socially relevant individual differences De Raad et al. (2016). The pioneering attempts date back to the early 20th century Allport and Odbert (1936); Galton (1950), and the paradigm has then been extended to different psychological constructs in decades of development Cattell (1957); John et al. (1988); Aavik and Allik (2002); Ponizovskiy et al. (2020); Mai and Church (2023); Garrashi et al. (2024).
Most of these works, taking values as an example, involve the following steps: 1) compilation of value-descriptive terms, mostly from dictionaries and thesauruses; 2) refinement and reduction to eliminate redundant, ambiguous, and uncommon terms; 3) value measurement, through collecting self-reports and peer ratings, to identify underlying correlations between the value descriptors; and 4) uncovering the hidden value factors or dimensions through statistical methods like principal component analysis or factor analysis.
In this work, we aim to harness the extensive knowledge and semantic understanding of LLMs to address the limitations of traditional psycho-lexical approaches. These limitations include: 1) extensive manual labor required to compile and refine the lexicons, 2) insufficient coverage of values of different linguistic forms, 3) lack of criteria for prioritizing values lexicons when filtering, 4) the responses bias and inscalability of self-report value measurement, and 5) the inadaptability to changing values or specific contexts.
@@ 2.3 From Human Values to LLM Values
The rise of LLMs also brings the critical need to evaluate, understand, and align their values (Ye et al., 2025a). Extensive works focus on evaluating the values of LLMs using traditional self-report tools Li et al. (2022, 2024c); Huang et al. (2024a); Safdari et al. (2023a, b), static customized inventories Ren et al. (2024); Meadows et al. (2024); Jiang et al. (2024a), dynamically generated inventories Jiang et al. (2024b), or directly from the model’s free-form outputs Ye et al. (2025b); Yao et al. (2024b, a, 2025). Other works attempt to understand the LLM values in aspects like the consistency of shown values Rozen et al. (2024); Moore et al. (2024); Röttger et al. (2024), the ability to reason about human values Ganesan et al. (2023); Sorensen et al. (2024a); Jiang et al. (2024d); Strachan et al. (2024), and the ability to express or role-play human values Jiang et al. (2024c); Zhang et al. (2023); Kang et al. (2024); Zhang et al. (2025); Li et al. (2024a); Huang et al. (2024b). Another line of research aligns LLM values with human values, setting the alignment goal as specific risk metrics Lin et al. (2024); Gunjal et al. (2024), human demonstrations Dubois et al. (2024); Köpf et al. (2024); Taori et al. (2023), implicit preference modeling Ouyang et al. (2022); Rafailov et al. (2024); Zhong et al. (2024), or intrinsic values Bai et al. (2022a); Yao et al. (2024a); Bai et al. (2022c). Among different alignment goals, intrinsic values, as transituational decision-guiding principles, demonstrate unique advantages Yao et al. (2023). However, prior works are mostly based on Schwartz’s value system, which is established and validated for humans and may not necessarily capture the unique LLM psychology.
One preliminary attempt Biedma et al. (2024) at constructing LLM value systems faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is susceptible to response bias and lacks comprehensive coverage; 2) the derivation of value correlations lacks theoretical or empirical validity; and 3) the resulting value system has yet to be evaluated in terms of its reliability, validity, or utility. Appendix C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
Generative Psycho-Lexical Approach (GPLA), as our foundation for constructing the LLM value system, is illustrated schematically in Fig. 1 and algorithmically in Algorithm 1. GPLA adopts an agentic framework, comprising three LLM agents and four sequential steps, to systematically collect, filter, measure, and structure the LLM value system.
@@
In line with the established psycho-lexical approach De Raad et al. (2016), we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs from ValueBench Ren et al. (2024), GPV Ye et al. (2025b), BeaverTails Ji et al. (2024), and ValueLex Biedma et al. (2024). These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in § A.1.
@@
Then, we extract perceptions from the collected corpus using the Perception Parser . Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools Ye et al. (2025b). These perceptions are then mapped to underlying values through the Value Generator , with their frequency of occurrence recorded. We instantiate following GPV Ye et al. (2025b) and using Kaleido model Sorensen et al. (2024a). In § B.1, we confirm the comprehensive coverage of the generated values.
@@
To ensure the value lexicons are concise and representative, we filter out duplicates using Rouge scores and embedding similarity following the validated setup in Sorensen et al. (2024a). When two values exhibit significant semantic overlap, we retain the one more frequently observed, as higher frequency signifies greater priority Ponizovskiy et al. (2020). This process results in value lexicons that minimize redundancy while preserving semantic coverage.
@@
Structuring the LLM value system requires identifying correlations between values based on LLM value measurements. To this end, we adopt GPV Ye et al. (2025b), instantiated with and , to measure the value orientations of LLM subjects on values . GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein (detailed in § A.3). We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (listed in § A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system is then constructed through Principal Component Analysis (PCA) Ponizovskiy et al. (2020), which uncovers the underlying value factors and hierarchical relationships (loadings of on ).
4 Benchmarking LLM Value Systems
The power of a value system lies in its validity and utility Schwartz (2012). This work introduces, for the first time, three benchmarking tasks for LLM value systems, grounded in both psychological theories and AI priorities: 1) Confirmatory Factor Analysis, which evaluates their structure validity given LLM value measurements; 2) LLM Safety Prediction, which assesses their predictive validity for LLM safety; and 3) LLM Value Alignment, which examines their representation power in aligning LLMs with human values.
4.1 Confirmatory Factor Analysis for Evaluating Structure Validity
Confirmatory Factor Analysis (CFA) is a statistical technique used to test whether a set of observed variables (atomic values) accurately represents and loads onto a smaller number of underlying latent factors (high-level values) Schwartz and Boehnke (2004). Mathematically, the CFA model is represented as:
4.2 LLM Safety Prediction for Evaluating Predictive Validity
Predictive validity refers to how well measurement results based on a particular value system can anticipate future decisions, actions, or outcomes Bardi and Schwartz (2003). In the context of human values, evaluating predictive validity often examines the correlations between values and other psychological constructs, such as personality traits and attitudes. For LLMs, we propose to evaluate the predictive validity by examining the relationship between LLM values and LLM safety, arguably one of the most critical concerns and "psychological constructs" for LLM deployment.
To this end, we draw prompts and safety scores from the established safety benchmark SALAD-Bench Li et al. (2024b). We administer the prompts to LLMs, collect their responses, and measure the revealed values using GPV Ye et al. (2025b). We follow the linear probing protocol Chen et al. (2020) to evaluate the predictive validity of the value representations under different value systems. Using the Bradley-Terry model Bradley and Terry (1952), we train a linear classifier with pairwise cross-entropy loss to predict the relative safety of LLM pairs:
4.3 LLM Value Alignment for Evaluating Representation Power
Values are cognitive representations of motivational goals Schwartz (2012). A well-constructed value system should be able to fully represent the broad motivations behind LLM outputs, therefore enabling effective transituational value alignment.
Our evaluation builds on the BaseAlign algorithm Yao et al. (2024a), originally designed to align LLM outputs with Schwartz’s basic human values. This approach employs a target human value vector, either heuristically customized or derived from human survey data. We extend BaseAlign to arbitrary value systems by distilling the target vector from human preference:
Here, is the target value vector for values under evaluation. It is distilled from dataset Ouyang et al. (2022) using open-vocabulary value evaluator Ye et al. (2025b). contains triplets of prompt , winning response , and losing response . The distillation objective is to minimize the L1 distance between the target value and the values of winning responses, while maximizing the distance with those of losing responses. Given that the objective function is piecewise linear, the optimal solution can be derived by iteratively examining each value dimension and the corresponding value measurements :
Next, we adopt PPO Schulman et al. (2017) to align the LLM with the target value under Yao et al. (2024a):
where is the LLM policy parameters, is the dataset of prompts Dai et al. (2023), and is the parameterized LLM response distribution. The representation power of is evaluated by the helpfulness and harmlessness of the aligned LLMs Bai et al. (2022b); Yao et al. (2024a).
[…]
5 Results
This section presents the results of our experiments, including the proposed LLM value system, the analysis of the system, and the benchmarking of our system against the well-established Schwartz’s values Schwartz (2012).
| Factor | Value | Loading |
| Social Responsibility (, CI: 0.952–0.961) | ||
| Equity | 0.890 | |
| Empathy | 0.885 | |
| Teamwork | 0.872 | |
| Equality | 0.827 | |
| Unity | 0.825 | |
| Risk-Taking (, CI: 0.910–0.928) | ||
| Challenge | 0.812 | |
| Boldness | 0.804 | |
| Adventure | 0.798 | |
| Change | 0.730 | |
| Thrill-seeking | 0.728 | |
| Rule-Following (, CI: 0.824–0.859) | ||
| Realism | 0.708 | |
| Order | 0.694 | |
| Responsibility | 0.682 | |
| Prudence | 0.676 | |
| Efficiency | 0.669 | |
| Self-Competence (, CI: 0.732–0.787) | ||
| Confidence | 0.601 | |
| Impact | 0.527 | |
| Proactivity | 0.460 | |
| Achievement | 0.455 | |
| Recognition | 0.455 | |
| Rationality (, CI: 0.686–0.754) | ||
| Objectivity | 0.705 | |
| Evidence-based | 0.649 | |
| Logic | 0.618 | |
| Neutrality | 0.522 | |
5.1 Proposed Value System
Fig. 3 visualizes the value system constructed by GPLA. Table 1 gathers its partial factor loadings, Cronbach’s Alpha (), and confidence intervals (CI). The factor loadings indicate the strength of the relationship between each factor and its atomic values, with higher loadings suggesting more contribution to the factor. Cronbach’s Alpha measures the internal consistency of each factor, with higher values indicating greater reliability. Some atomic values are removed to ensure clear loading patterns and desirable factor reliability Aavik and Allik (2002). After that, all our factors reach the standard psychometric threshold of 0.7, indicating strong internal consistency Taber (2018).
Analyzing factor loadings reveals the underlying structure of five key factors: Social Responsibility emphasizes fairness, inclusivity, and societal well-being; Risk-Taking reflects openness to change and embracing uncertainty; Rule-Following highlights order, discipline, and pragmatism; Self-Competence focuses on personal growth and achievement; and Rationality prioritizes logical, evidence-based decision-making grounded in objectivity and impartiality. Full factor loadings and further analysis are available in § B.3.
5.2 Analyzing Value System
Value Correlation Analysis.
Fig. 3 (Top) presents the correlations between the factors in our value system. Similar to Schwartz’s theory of basic human values Schwartz (2012), LLM values also exhibit both compatible and opposing relationships. Notably, social responsibility, rule-following, and rationality show positive correlations with one another, while all of them are negatively correlated with risk-taking.
Circumplex Analysis.
Circumplex analysis is a statistical method that examines whether the underlying structure of variables aligns with a circumplex pattern, and, if so, the positions of variables on a circle. The stronger the correlation between variables, the shorter their distance on the circumference. We conduct circumplex analysis based on the correlations between factors. Fig. 3 (Bottom) illustrates the analysis results based on Browne’s circular stochastic process model Browne (1992); Grassi et al. (2010). The compatible values are closer on the circle (e.g., Social Responsibility and Rule-Following) while opposing values are positioned diagonally (e.g., Risk-Taking and Rule-Following). The results verify the presence of a circumplex structure in the value system.
Consistency Across Datasets.
To evaluate the consistency and robustness of our multivariate system structure (i.e., the 5-dimensional relationship), we measure LLM values using two distinct prompt datasets: the psychometric prompts from GPV Ye et al. (2025b) and the red-teaming prompts from SALAD-Bench Li et al. (2024b), which feature distinct prompt distributions. Their measurements yield an average intra-LLM correlation of 0.87; here we use intra-LLM correlations because the relative value hierarchy within an LLM is more important than their absolute measurements. This result indicates a high level of consistency in the value structure across prompt distributions. We also find a high correlation (0.73) between intra-LLM value consistency and LLM safety scores Li et al. (2024b). It suggests that LLMs with higher value consistency tend to be safer. Complete results are available in § B.4.
5.3 Comparing Value Systems
We benchmark the proposed value system against Schwartz’s value system Schwartz (2012), the most established framework for human values and commonly used in LLM value studies.
Confirmatory Factor Analysis.
We follow the standard validation procedures of CFA Schwartz and Boehnke (2004) to evaluate the structure validity of different value systems. For our value system, we construct it using half of the measurement data and bootstrap the data to ensure its sufficiency (#data points #variables). The other half of the data is held out for CFA. For Schwartz’s value systems, we map the observed value variables (the atomic values in our system) to its four high-level values or ten low-level values, according to the semantic relevance Schwartz and Boehnke (2004) measured by an embedding model OpenAI (2024); all data is used for CFA. Table 2 displays the CFA results. Our value system demonstrates a better fit for the data, capturing the underlying values behind LLM generations.
| Value system | Schwartz (H) | Schwartz (L) | Ours |
| #Values | 4 | 10 | 5 |
| CFI | 0.56 | 0.23 | 0.68 |
| GFI | 0.52 | 0.22 | 0.65 |
| RMSEA | 0.10 | 0.11 | 0.12 |
| AIC | 340 | 324 | 265 |
| BIC | 1484 | 1464 | 1145 |
LLM Safety Prediction.
We follow the setup in (Ye et al., 2025b, Section 5.2) to predict LLM safety based on their value orientations. Table 3 presents the prediction accuracy under different value systems. The higher accuracy of our value system indicates its superior predictive validity. In addition, according to the parameters of well-trained linear classifiers, we find that Social Responsibility, Rule-Following, and Rationality enhance safety, whereas Risk-taking and Self-Competence undermine it; see § B.6 for full results.
| Value system | Acc (%) |
| Schwartz (H) | 81 15 |
| Schwartz (L) | 74 16 |
| Ours | 87 9 |
LLM Value Alignment.
We follow the experimental setup in (Yao et al., 2024a, Section 6.2) to perform LLM value alignment. Different value systems are respectively used to represent LLM outputs and desired human values. We employ GPV Ye et al. (2025b) as an open-vocabulary value evaluator for all value systems, but also include the original results of Yao et al. (2024a) using a Schwartz-specific evaluator for comparison. Table 4 shows the alignment performance of different value systems. Alignment under our value system converges to the lowest harmlessness and the highest helpfulness, establishing its superior representation power. Full experimental details are available in § A.4.
| Value system | Harmlessness | Helpfulness |
| Schwartz* | -1.52 | 2.15 |
| Schwartz | -1.40 | 2.13 |
| Ours | -1.26 | 2.16 |
6 Discussion
6.1 Validation of LLM Capabilities